# Decoding Susceptibility: Modeling Misbelief to Misinformation Through a Computational Approach

### v2

@@ Abstract
Susceptibility to misinformation describes the degree of belief in unverifiable claims, a latent aspect of individuals’ mental processes that is not observable. Existing susceptibility studies heavily rely on self-reported beliefs, which can be subject to bias, expensive to collect, and challenging to scale for downstream applications. To address these limitations, in this work, we propose a computational approach to model users’ latent susceptibility levels. As shown in previous research, susceptibility is influenced by various factors (e.g., demographic factors, political ideology), and directly influences people’s reposting behavior on social media. To represent the underlying mental process, our susceptibility modeling incorporates these factors as inputs, guided by the supervision of people’s sharing behavior. Using COVID-19 as a testbed domain, our experiments demonstrate a significant alignment between the susceptibility scores estimated by our computational modeling and human judgments, confirming the effectiveness of this latent modeling approach. Furthermore, we apply our model to annotate susceptibility scores on a large-scale dataset and analyze the relationships between susceptibility with various factors. Our analysis reveals that political leanings and psychological factors exhibit varying degrees of association with susceptibility to COVID-19 misinformation. 111We will release all the code used in our paper, along with our trained model and all collected data.
@@ 1 Introduction
People’s mental processes, which are unobservable and influenced by various factors, directly affect several externalized behaviors, such as reposting on social media Mitchell et al. (2019); Brady et al. (2020); Islam et al. (2020); Altay et al. (2022). Building on these prior works, we propose a computational method to model individuals’ unobservable susceptibility levels, utilizing user-related factors and the supervision of their observable social media sharing behaviors. More specifically, our latent susceptibility modeling process uses a user’s historical tweets as input to predict their sharing behavior for specific misinformation. We perform multi-task learning to simultaneously learn to classify whether a user would share a post, and rank susceptibility scores among similar and dissimilar users when the same content is seen. This computational modeling method unlocks the scales of misinformation-related studies and provides a novel perspective to reveal users’ belief patterns.
In this paper, we focus our experiments on COVID-19 misinformation, and our evaluations demonstrate that the estimations from our model are highly aligned with human judgment when assessed through a susceptibility comparison task. We further employ our model to annotate susceptibility levels on a large-scale dataset. Building upon this extensive susceptibility labeling, we then conduct a set analysis to examine how various social factors relate to susceptibility. Our analysis reveals that psychological factors, professional fields, and political leanings are associated with susceptibility to varying degrees. Notably, this analysis, grounded in our inferred susceptibility scores, corroborates the findings of previous studies based on self-reported beliefs—for example, confirming that stronger analytical thinking is an indicator of lower susceptibility. Moreover, the results of our analysis show the potential to extend findings in the existing literature. For example, we demonstrate that the distribution of COVID-19 misinformation susceptibility in the U.S. exhibits a certain degree of correlation with political leanings. To sum up, our contributions are:
We propose a computational model to infer people’s susceptibility to misinformation by leveraging observable sharing activities to model unobservable latent susceptibility.
Our evaluation demonstrates that our model effectively captures unobservable susceptibility levels, and its predictions on the susceptibility comparison task closely align with human judgment.
We conduct a large-scale analysis to identify the underlying factors influencing susceptibility among a diverse user population, spanning various professional and geographical communities, thereby offering valuable insights for related social science research.
@@ 2 Related Work
The common practice to measure susceptibility is to collect self-reported absolute or relative agreement or disagreement with (or perceived accuracy, credibility, reliability, or validity of) one or more claims verified to be false from a group of individuals Roozenbeek et al. (2020); Escolà-Gascón et al. (2021); Rosenzweig et al. (2021); Nan et al. (2022). A small number of previous studies indirectly assess the susceptibility by its impact, however, they can only capture behaviors rather than people’s beliefs Loomba et al. (2021). Cheng et al. (2021) defines a heuristic susceptibility score as the radio of misinformation posts out of all user’s posts, which unrealistically simplifies the definition of susceptibility. Instead of using expensive and limited self-reported beliefs, we propose a computational model to estimate susceptibility at scale.
@@
Relying on the manually collected susceptibility annotation, previous research investigates the psychological, demographic, and more factors that contribute to users’ susceptibility Bringula et al. (2021); van der Linden (2022). These factors include emotion Sharma et al. (2022) (e.g. anger and anxiety; Weeks, 2015), analytic thinking hui Li et al. (2022), partisan bias Roozenbeek et al. (2022a), source credibility Traberg and van der Linden (2022), and repetition Foster et al. (2012). Many theories have been proposed about the reason behind suscetibility Scherer et al. (2020), including limited knowledge acquiring and literacies capabilities Brashier and Schacter (2020), strong preexisting beliefs Lewandowsky and Ecker (2012), neglecting to sufficiently reflect about the truth Pennycook and Rand (2017) or overconfidence Salovich et al. (2020).
A better understanding of the phenomenon and mechanism of susceptibility can facilitate various downstream applications. These include analyzing the spread of bots Himelein-Wachowiak et al. (2021), revealing community properties in information pathways Taylor et al. (2023); Ma et al. (2023), combating misinformation by emphasizing publisher Dias et al. (2020) and prebunking interventions based on inoculation Roozenbeek et al. ; Roozenbeek et al. (2022b). However, the absence of a computational modeling framework significantly limits the scale of current susceptibility research.
@@ 3 Computational Susceptibility Modeling
Misinformation is characterized as information that is false, inaccurate, or misleading, which could be created deliberately or accidentally Pennycook and Rand (2021). The susceptibility to misinformation represents the belief in misinformation and related constructs, including discernment between true and false claims and the extent to which exposure to misinformation misleads subsequent decisions Nan et al. (2022). Believing a piece of misinformation can influence various outward behaviors, such as sharing actions. Previous studies, like the inattention or “classical reasoning” account, contend that people are committed to sharing accurate information, but the unique context of social media disrupts their capacity to critically assess the accuracy of news Pennycook and Rand (2021); van der Linden (2022). These studies suggest that people are more likely to share things they genuinely believe Altay et al. (2022).
Inspired by the existing studies indicating that believing is an essential driver for dissemination, we propose to model the latent susceptibility, which reflects people’s beliefs, with supervision from the sharing behaviors, while considering characteristics of the sharing content and user representation. We introduce a computational model to infer a user’s unobservable susceptibility level based on their historical activities (§3.1) as shown in Fig. 1. We train the model with signals from the observable sharing behaviors on social media(§3.2), utilizing the constructed approximate contrastive user-post pairs as the training data (§4).
@@ 3.1 Modeling Unobservable Susceptibility
We compute the susceptibility of user when a particular piece of misinformation is perceived (i.e. ). This allows us to account for the fact that an individual’s susceptibility can vary across different content, influenced by factors such as topics and linguistic styles. By focusing on the susceptibility to specific pieces of misinformation, we aim to create a more nuanced, fine-grained, and accurate representation of how users interact with and react to different misinformation.
@@
As a component of the computational model, we use SBERT Reimers and Gurevych (2019), which is developed upon RoBERTa-large Liu et al. (2019), to compute the embedding vector to represent the semantic information contained in the misinformation and user historical posts. We consider the misinformation post as a sentence and produce its representation with SBERT. For the user embedding, we calculate the average of sentence representations for the user’s recent original posts. More specifically, for every user-post pair , we gather the historical posts written by user within a 10-day window preceding the creation time of the misinformation post , to learn a representation of user at that specific time.222We chose the 10-day timeframe because it provides a substantial amount of data to represent a user and is also recent enough to capture their dynamics.
@@ 4 Dataset and Experiment Setup
We have chosen Twitter as our data source because it hosts an extensive and diverse collection of users, the accessibility of its data, and its popularity for computational social science research Zhang and Counts (2015). Furthermore, Twitter offers users the freedom for free-text personal and emotional expression, with crucial metadata, including timestamps and location data, which are useful for our subsequent analytical endeavors.
@@
Regarding negative examples, we define them as pairs where user viewed but did not retweet misinformation post . However, obtaining these negative examples poses a considerable challenge, because the Twitter API does not provide information on the “being viewed” activities of a specific tweet. To address this issue, we construct potential negative users who are highly likely to have viewed a particular post but did not repost it, following these heuristics: 1) should be a follower of the author of the misinformation post , 2) should not retweet , and 3) was active on Twitter within 10 days before and 2 days after the timestamp of .
In the end, we collected 3,811 positive examples and 3,847 negative examples, resulting in a dataset consisting of a total of 7,658 user-post pairs. We divide the dataset into three subsets with an 80% - 10% - 10% split for train, validation, and test purposes, respectively. The detailed statistics of the collected data are illustrated in Tab. 1. We provide the training details of our model in Appendix A.
@@ 5.1 Validation with Human Judgement
Due to the abstract nature of susceptibility and the absence of concrete ground truth, we encounter challenges in directly assessing our susceptibility modeling. As a result, we tend to human evaluations to validate the effectiveness of our modeled susceptibility. Given the inherent subjectivity in the concept of susceptibility, and to mitigate potential issues arising from variations in individual evaluation scales, we opt not to request humans to annotate a user’s susceptibility directly. Instead, we structure the human evaluation as presenting human evaluators with pairs of users along with their historical tweets and requesting them to determine which user appears more susceptible to overall COVID-19 misinformation. We provide more details regarding the human judgment framework and the utilized interface in Appendix B.
Subsequently, we compared the predictions made by our model with the human-annotated predictions. To obtain predictions from our model, we compute each user’s susceptibility to overall COVID-19 misinformation by averaging their susceptibility scores to each COVID-19 misinformation tweet in our dataset. As presented in Tab. 2, our model achieves an agreement of 72.90% with human predictions, indicating a solid alignment with the annotations provided by human evaluators. Additionally, we consider a baseline that directly calculates susceptibility scores as the cosine similarity between the user and misinformation tweet embeddings. Compared to this baseline, our susceptibility modeling brings a 9.35% improvement. Moreover, we conduct a comparison with ChatGPT by providing it with instructions based on the task description of the susceptibility level comparison setting in a zero-shot manner (more details are in Appendix D). We notice that our model even outperforms predictions made by ChatGPT, despite ChatGPT being a significantly larger model than ours. These results of the human judgment validate the effectiveness of our proposed susceptibility modeling, showcasing its capability to reliably estimate user susceptibility to COVID-19 misinformation.
@@ 6 Analysis
Furthermore, we demonstrate in this section the potential of our proposed computational susceptibility modeling in expanding the scope of susceptibility-related research. Our analysis not only aligns with the findings of previous survey-based studies (§6.1) but also goes a step further by extending and enriching their conclusions (§6.2).
@@ 6.2 Community Differences
To perform a reliable analysis among different communities, a large-scale user dataset is needed. To address this requirement, we sample 100,000 users across the world from the existing COVID-19 Tweet Dataset Taylor et al. (2023) which contains all COVID-19-related tweets for a certain time. 444Besides, we make sure each sampled user has posted more than 100 historical tweets between January 2020 and April 2021. For each user, we utilize the Twitter API to gather their user descriptions and location information, after which we extract and categorize their occupations from their self-reported descriptions with ChatGPT in a zero-shot manner. Furthermore, for better interpretability, we normalize the resulting susceptibility scores within the range of -100 to 100 using Min-Max normalization. We define -100 to indicate that the individual holds the most resistance to misinformation, while 100 means the individual is easiest to believe in misinformation when encountered. To obtain an aggregated susceptibility score for a community, we calculate the mean of individual susceptibility scores for all users within that community.
@@
The results are presented in Tab. 4. It is worth noting that professions closely associated with S&T, H&M, and Edu tend to exhibit lower susceptibility to COVID-19 misinformation. In contrast, occupations within the A&M area demonstrate comparatively higher susceptibility, possibly because of their greater exposure to misinformation and stronger emotional reactions. These findings align with our expectations and reinforce the previous conclusions that expertise and knowledge in relevant fields serve as protective factors against misinformation Nan et al. (2022)555We notice that Twitter users who don’t declare their occupation in their user description (N/A) exhibit a higher susceptibility to COVID misinformation. This may be because those who are willing to declare their profession are often public figures who care more about their reputation..
@@
We further investigate the geographical distribution of susceptibility to COVID-19 misinformation, specifically focusing on different U.S. states.666Given the imbalance in the number of users from different U.S. states, we calculate average susceptibility scores for each state with Bayesian smoothing. We use the overall mean and overall standard deviation as priors, and the more users in the state, the less the overall mean will affect that state’s score. This analysis enables us to explore the influence of political ideology associated with different U.S. states Gelman (2008) on susceptibility to misinformation. Out of the 100,000 users sampled from around the world, 25,653 users are from U.S. states with more than 200 users for each state. As shown in Fig. 3, the distribution of susceptibility levels estimated by our computational modeling is imbalanced across U.S. states and demonstrates a certain degree of correlation with political leanings. In general, states known to have a more conservative population tend to have relatively higher susceptibility scores, while states that are considered more liberal have lower scores. The average susceptibility score for users in blue or red states is -3.66 and -2.82 respectively.777Red and blue states are determined by the 2020 presidential election results, with red states leaning Republican and blue states leaning Democratic. We observe that 60% of the ten states with the highest susceptibility scores are red states, and 90% of the ten states with the lowest susceptibility scores are blue states. This trend corresponds with the conclusion observed in various previous studies, where political ideology influences people’s perspectives on scientific information McCright et al. (2013); Baptista et al. (2021); Imhoff et al. (2022). However, it is important to acknowledge the limitations of our analysis, as it solely reflects the estimated susceptibility distribution of the sampled users within each state. In Appendix E, we present the average susceptibility scores calculated based on our sampled users for each U.S. state, along with the corresponding number of users.

### v3

@@ Abstract
Susceptibility to misinformation describes the degree of belief in unverifiable claims, a latent aspect of individuals’ mental processes that is not observable. Existing susceptibility studies heavily rely on self-reported beliefs, which can be subject to bias, expensive to collect, and challenging to scale for downstream applications. To address these limitations, in this work, we propose a computational approach to efficiently model users’ latent susceptibility levels. As shown in previous work, susceptibility is influenced by various factors (e.g., demographic factors, political ideology), and directly influences people’s reposting behavior on social media. To represent the underlying mental process, our susceptibility modeling incorporates these factors as inputs, guided by the supervision of people’s sharing behavior. Using COVID-19 as a testbed, our experiments demonstrate a significant alignment between the susceptibility scores estimated by our computational modeling and human judgments, confirming the effectiveness of this latent modeling approach. Furthermore, we apply our model to annotate susceptibility scores on a large-scale dataset and analyze the relationships between susceptibility with various factors. Our analysis reveals that political leanings and other psychological factors exhibit varying degrees of association with susceptibility to COVID-19 misinformation, and shows that susceptibility is unevenly distributed across different professional and geographical backgrounds.111We will release all the code used in our paper, along with our trained model and all collected data.
Decoding Susceptibility: Modeling Misbelief to Misinformation Through a Computational Approach
Yanchen Liu
Mingyu Derek Ma
Wenna Qin
Azure Zhou
Jiaao Chen
Weiyan Shi
Wei Wang
Diyi Yang
Harvard University
Stanford University
UCLA
Georgia Institute of Technology
yanchenliu@g.harvard.edu, {wennaqin, amysz, weiyans, diyiy}@cs.stanford.edu,
{ma, weiwang}@cs.ucla.edu, jiaaochen@gatech.edu
@@ 1 Introduction
People’s mental processes, which are unobservable and influenced by various factors, directly affect several externalized behaviors, such as reposting on social media Mitchell et al. (2019); Brady et al. (2020); Islam et al. (2020); Altay et al. (2022). Building on these prior works, we propose a computational method to efficiently model individuals’ unobservable susceptibility levels only based on their observable social media posting and sharing behaviors. We represent users based on their historical posts and perform multi-task learning to simultaneously learn to classify whether a user would share a post, as well as to rank susceptibility scores among similar and dissimilar users when the same content is seen. This computational modeling method unlocks the scales of misinformation-related studies and provides a novel perspective to reveal users’ belief patterns.
In this paper, we focus our experiments on COVID-19 misinformation, and our evaluations demonstrate that the estimations from our model are highly aligned with human judgment when assessed through a susceptibility comparison task. The correlation study between estimated and human-annotated susceptibility verifies the effectiveness of the indirect susceptibility modeling method. To further illustrate the significance of our work, we employ our model to annotate susceptibility levels on a large-scale dataset. Building upon this extensive susceptibility labeling, we then conduct a set analysis to examine how various factors relate to susceptibility. Our analysis reveals that psychological factors, professional fields, and political leanings are associated with susceptibility to varying degrees. Notably, this large-scale analysis enabled by our computational susceptibility modeling corroborates the findings of previous studies based on self-reported beliefs, e.g. confirming that stronger analytical thinking is an indicator of lower susceptibility. Moreover, the results of our analysis show the potential to extend findings in the existing literature. For example, we demonstrate that the distribution of COVID-19 misinformation susceptibility in the U.S. exhibits a certain degree of correlation with political leanings.
@@ 2 Related Work
The common practice to measure susceptibility is to collect self-reported absolute or relative agreement or disagreement with (or perceived accuracy, credibility, reliability, or validity of) one or more claims verified to be false from a group of individuals Roozenbeek et al. (2020); Escolà-Gascón et al. (2021); Rosenzweig et al. (2021); Nan et al. (2022). A small number of previous studies indirectly assess the susceptibility by its impact, however, they can only capture behaviors rather than people’s beliefs Loomba et al. (2021). Cheng et al. (2021) defines a heuristic susceptibility score as the ratio of misinformation posts out of all user’s posts, which unrealistically simplifies the definition of susceptibility. Instead of using expensive and limited self-reported beliefs, we propose a computational model to estimate susceptibility at scale.
@@
Relying on the manually collected susceptibility annotation, previous research investigates the psychological, demographic, and more factors that contribute to users’ susceptibility Bringula et al. (2021); van der Linden (2022). These factors include emotion Sharma et al. (2022) (e.g. anger and anxiety; Weeks, 2015), analytic thinking hui Li et al. (2022), partisan bias Roozenbeek et al. (2022a), source credibility Traberg and van der Linden (2022), and repetition Foster et al. (2012). Many theories have been proposed about the reason behind suscetibility Scherer et al. (2020), including limited knowledge acquiring and literacies capabilities Brashier and Schacter (2020), strong preexisting beliefs Lewandowsky and Ecker (2012), neglecting to sufficiently reflect about the truth Pennycook and Rand (2017) or overconfidence Salovich et al. (2020).A better understanding of the phenomenon and mechanism of susceptibility can facilitate various downstream applications. These include analyzing the spread of bots Himelein-Wachowiak et al. (2021), revealing community properties in information pathways Taylor et al. (2023); Ma et al. (2023), combating misinformation by emphasizing publisher Dias et al. (2020) and prebunking interventions based on inoculation Roozenbeek et al. ; Roozenbeek et al. (2022b). However, the absence of a computational modeling framework significantly limits the scale of current susceptibility research.
@@ 3 Computational Susceptibility Modeling
Misinformation is characterized as information that is false, inaccurate, or misleading, which could be created deliberately or accidentally Pennycook and Rand (2021). The susceptibility to misinformation represents the belief in misinformation and related constructs, including discernment between true and false claims and the extent to which exposure to misinformation misleads subsequent decisions Nan et al. (2022). Previous research on susceptibility and misinformation mainly relied on self-reported beliefs collected using surveys or questionnaires - they suffered from problems like being subject to bias, expensive to collect, and challenging to reproduce and scale up.
Existing studies indicating that believing a piece of misinformation can influence various outward behaviors, such as sharing actions. For example, previous studies of the inattention or “classical reasoning” account contend that people are committed to sharing accurate information, but the unique context of social media disrupts their capacity to critically assess the accuracy of news Pennycook and Rand (2021); van der Linden (2022). These studies suggest that people are more likely to share things they genuinely believe Altay et al. (2022). Inspired by this observation, we propose to model user’s unobservable susceptibility only based on their historical posting and sharing behaviors, which are the most available and the easiest collectable data from social media (§3.1) as shown in Fig. 1. Therefore, our proposed framework can efficiently infer users’ susceptibility levels to misinformation on a large scale, demonstrating the potential to expand the scope of previous misinformation-related research.
Furthermore, because social media users utilize posts to express their personal and inner thoughts, they reveal information about their characteristics through their posts. Therefore, our proposed susceptibility modeling can incorporate users’ informative hidden factors, such as personality traits, analytical thinking, and emotion, to infer a user’s susceptibility to misinformation. These additional pieces of information are otherwise very difficult to directly collect on social media.
@@ 3.1 Modeling Unobservable Susceptibility
In our work, we consider the susceptibility of user when a particular piece of misinformation is perceived (i.e. ). This allows us to account for the fact that an individual’s susceptibility can vary across different content, influenced by factors such as topics and linguistic styles. By focusing on the susceptibility to specific pieces of misinformation, we aim to create a more nuanced, fine-grained, and accurate representation of how users interact with and react to different misinformation.
@@
We induce user and post embeddings to reflect hidden factors of the user personality traits and content of the post. As a component of the computational model, we use SBERT Reimers and Gurevych (2019), which is developed upon RoBERTa-large Liu et al. (2019), to compute the embedding vector to represent the information contained in the misinformation and user historical posts. We consider the misinformation post as a sentence and produce its representation with SBERT. For the user embedding, we calculate the average of sentence representations for the user’s recent original posts. More specifically, for every user-post pair , we gather the historical posts written by user within a 10-day window preceding the creation time of the misinformation post , to learn a representation of user at that specific time.222We chose the 10-day timeframe because it provides a substantial amount of data to represent a user and is also recent enough to capture their dynamics.
@@
Scale and Interpretation of Susceptibility Score
Furthermore, for better interpretability, we normalize the resulting susceptibility scores within the range of -100 to 100 using Min-Max normalization. We define -100 to indicate that the individual holds the most resistance to misinformation, while 100 means the individual is easiest to believe in misinformation when encountered.
@@ 4 Dataset and Experiment Setup
We have chosen Twitter as our data source because it hosts a diverse collection of users and allows for free-text personal and emotional expression. Furthermore, Twitter provides crucial metadata, including timestamps and location data, which are useful for our subsequent analysis.
@@
Regarding negative examples, we define them as pairs where user viewed but did not retweet misinfo post . However, obtaining these negative examples poses a considerable challenge, because the Twitter API does not provide information on the “being viewed” activities of a specific tweet. To address this issue, we construct potential negative users who are highly likely to have viewed a particular post but did not repost it, following these heuristics: 1) should be a follower of the author of the misinformation post , 2) should not retweet , and 3) was active on Twitter within 10 days before and 2 days after the timestamp of .
In the end, we collected 3,811 positive examples and 3,847 negative examples, resulting in a dataset consisting of a total of 7,658 user-post pairs. We divide the dataset into three subsets with an 80% - 10% - 10% split for train, validation, and test purposes, respectively. The detailed statistics of the collected data are illustrated in Tab. 1. We provide the training details of our model in Appendix B.
@@ 5.1 Validation with Human Judgement
Due to the abstract nature of susceptibility and the absence of concrete ground truth, we encounter challenges in directly assessing our susceptibility modeling. As a result, we tend to human evaluations to validate the effectiveness of our modeled susceptibility. Given the inherent subjectivity in the concept of susceptibility, and to mitigate potential issues arising from variations in individual evaluation scales, we opt not to request humans to annotate a user’s susceptibility directly. Instead, we structure the human evaluation as presenting human evaluators with pairs of users along with their historical tweets and requesting them to determine which user appears more susceptible to overall COVID-19 misinformation. We provide more details regarding the human judgment framework and the utilized interface in Appendix C.
Subsequently, we compared the predictions made by our model with the human-annotated predictions. To obtain predictions from our model, we compute each user’s susceptibility to overall COVID-19 misinformation by averaging their susceptibility scores to each COVID-19 misinformation tweet in our dataset. As presented in Tab. 2, our model achieves an agreement of 72.90% with human predictions, indicating a solid alignment with the annotations provided by human evaluators. Additionally, we consider a baseline that directly calculates susceptibility scores as the cosine similarity between the user and misinformation tweet embeddings. Compared to this baseline, our susceptibility modeling brings a 9.35% improvement. Moreover, we conduct a comparison with ChatGPT by providing it with instructions based on the task description of the susceptibility level comparison setting in a zero-shot manner (more details are in Appendix E). We notice that our model even outperforms predictions made by ChatGPT, despite ChatGPT being a significantly larger model than ours. These results of the human judgment validate the effectiveness of our proposed susceptibility modeling, showcasing its capability to reliably estimate user susceptibility to COVID-19 misinformation.
@@ 6 Analysis
To further illustrate the significance of our work for the Computational Social Science community in susceptibility and misinformation research, we conducted a large-scale analysis on our collected large Twitter datasets and analyzed the correlation between user’s susceptibility and their psychological factors (§6.1), professional backgrounds (§6.2), and geographical distribution (§6.2). Our findings demonstrate that the large-scale analysis enabled by our proposed efficient susceptibility modeling not only corroborates the results of previous questionnaire-based studies, but also shows the potential of further extending the scope of research on susceptibility and misinformation.
@@ 6.2 Community Differences
To perform a reliable analysis among different communities, a large-scale user dataset is needed. To address this requirement, we sample 100,000 users across the world from the existing COVID-19 Tweet Dataset Taylor et al. (2023) which contains all COVID-19-related tweets for a certain time444Besides, we make sure each sampled user has posted more than 100 historical tweets between January 2020 and April 2021. For each user, we utilize the Twitter API to gather their user descriptions and location information, after which we extract and categorize their occupations from their self-reported descriptions with ChatGPT in a zero-shot manner.. To obtain an aggregated susceptibility score for a community, we calculate the mean of individual susceptibility scores for all users within that community.
@@
The results are presented in Tab. 4. It is worth noting that occupations within the A&M area demonstrate comparatively higher susceptibility, possibly because of their greater exposure to misinformation and stronger emotional reactions. In contrast, professions closely associated with S&T, F&B, H&M, S&P, and Edu tend to exhibit lower susceptibility to COVID-19 misinformation. These findings reinforce the previous conclusions that expertise and knowledge in relevant fields serve as protective factors against misinformation, especially for populations in the field of H&M Nan et al. (2022). Surprisingly, we notice that the S&T group is the most susceptible among the unsusceptible groups. For this, we have some speculations about the underlying reasons; for example, some people in the S&T community might have a higher level of skepticism towards traditional institutions and expertise, partly influenced by the culture that values disruptive innovation555We notice that Twitter users who don’t declare their occupation in their user description (N/A) exhibit a higher susceptibility to COVID misinformation. This may be because those who are willing to declare their profession are often public figures who care more about their reputation..
@@
We further investigate the geographical distribution of susceptibility to COVID-19 misinformation, specifically focusing on different U.S. states.666Given the imbalance in the number of users from different U.S. states, we calculate average susceptibility scores for each state with Bayesian smoothing. We use the overall mean and overall standard deviation as priors, and the more users in the state, the less the overall mean will affect that state’s score. This analysis enables us to explore the influence of political ideology associated with different U.S. states Gelman (2008) on susceptibility to misinformation. Out of the 100,000 users sampled from around the world, 25,653 users are from U.S. states with more than 200 users for each state. As shown in Fig. 3, the distribution of susceptibility levels estimated by our computational modeling is imbalanced across U.S. states and demonstrates a certain degree of correlation with political leanings. In general, states known to have a more conservative population tend to have relatively higher susceptibility scores, while states that are considered more liberal have lower scores. The average susceptibility score for users in blue or red states is -3.66 and -2.82 respectively.777Red and blue states are determined by the 2020 presidential election results, with red states leaning Republican and blue states leaning Democratic. We observe that 60% of the ten states with the highest susceptibility scores are red states, and 90% of the ten states with the lowest susceptibility scores are blue states. This trend corresponds with the conclusion observed in various previous studies, where political ideology influences people’s perspectives on scientific information McCright et al. (2013); Baptista et al. (2021); Imhoff et al. (2022). However, it is important to acknowledge the limitations of our analysis, as it solely reflects the estimated susceptibility distribution of the sampled users within each state. In Appendix F, we present the average susceptibility scores calculated based on our sampled users for each U.S. state, along with the corresponding number of users.